OpenStudy (anonymous):

I will give medal who answers this question. A farmer had twice as many chickens as ducks on his farm. After he sold 166 chickens and 29 ducks, he had half as many chickens as ducks left. How many chickens did the farmer have at first.

OpenStudy (anonymous):

please help

OpenStudy (eyust707):

Lets take these words, make them into equations then solve the equations.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank u for your time @eyust707

OpenStudy (eyust707):

"A farmer had twice as many chickens as ducks"

OpenStudy (anonymous):

yeah at first

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

no u didi wrong

OpenStudy (anonymous):

did

OpenStudy (eyust707):

C = 2 *D

OpenStudy (anonymous):

yep this is right

OpenStudy (anonymous):

next ?????

OpenStudy (anonymous):

can u do me a favor - please do it fast i have to go for an urgent work after 15 minutes

OpenStudy (anonymous):

thank u

OpenStudy (eyust707):

"After he sold 166 chickens and 29 ducks, he had half as many chickens as ducks left." Cf =C -166 Df = D-29 2 *Cf = Df

OpenStudy (anonymous):

what is Cf at the top

OpenStudy (eyust707):

Cf = final chickens Df = final ducks

OpenStudy (anonymous):

oh - next ???

OpenStudy (anonymous):

please try do do it as fast as u can

OpenStudy (anonymous):

thank u

OpenStudy (eyust707):

Substituting we get: 2*(C-166)= D-29 and we know from before that C = 2D 2*(2D-166)=D-29 simplify and solve for D 4D-332=D-29 4D-303=D 3D= 303 D=101

OpenStudy (eyust707):

C=2D C=2*101 C=202

OpenStudy (anonymous):

are we done did we get answer

OpenStudy (eyust707):

you tell me

OpenStudy (eyust707):

did you follow that?

OpenStudy (anonymous):

i think we got the answer

OpenStudy (anonymous):

which country do u live

OpenStudy (eyust707):

read it over, figure out what i did, try it for yourself, i think youll understand a lot better.

OpenStudy (anonymous):

you are from which country

OpenStudy (eyust707):

US

OpenStudy (anonymous):

which grade

OpenStudy (anonymous):

i am in 7 grade

OpenStudy (eyust707):

oh cool. Im in college

OpenStudy (anonymous):

oh u are good in math

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

are u still here

OpenStudy (anonymous):

let me give u medal sorry i forgot

OpenStudy (anonymous):

i gave u a medal

OpenStudy (anonymous):

thank u very much for your help

OpenStudy (anonymous):

u are the best person i have seen in world

OpenStudy (anonymous):

now can u go